Shift Summary for day shift, 08/12/99

Sandrine Laplace (laplace@hep.uchicago.edu)
Thu, 12 Aug 1999 15:44:27 -0500

Shift Summary for day shift, 08/12/99

On Shift: Sandrine (this is my last shift in KTeV !!!), Ashkan

Hours of beam delivered at all intensities: 6.2 hours
Hours of beam delivered at nominal intensity: 6.2 hours
Hours of E832 data (special runs not included): 5 hours
Hours of special runs (laser scans, muon runs...): 0 hours

Accounting of lost time: Beam quench and then other problems.
CsI HV problems.

========================================================================
Runs ended this shift:

14202 - BEAM_832 - 750k events
14203 - " - 13.7M events
14204 - " - bad run

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:

smooth 832 data taking until 12:24 when TeV quenches.
Then, after having verified CsI HVs, big problems happened with HV.
What is in the database is completely wrong.
At the end of the run, Julie and JasonH are still working on it.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14202 7:00 New crew arrives. Run in progress, trigger was
stopped (see previous shift summary) but fixed
by drt_fresh_start.

14203 7:10 starts run 14203.
Ran the 'usual' KFREND at the first spill.

8:30 Just finished to look at online histos:
-- L2 signal efficiency for eta+- is a little
bit low (0.998..).
-- HCCMON plot 'energy of all CsI blocks',
threshold seems too high (1.1 instead
of 1). Jiambo notified.

68 8:50 KTevana Error, ran KFREND.

9:00 channel 1023 has 4.68 stab.RMS and remains stable
at this value.
don't see comp.problems for channel 166.

9:35 Called MCR to ask them to keep SEM below 7e12.
They are doing some tests so it's gonna fluctuate
for
a short while.

104 9:40 KTevana Error, ran KFREND.

10:20 Channel 28 of HA shows a ped.shift of 100 counts.
Elizabeth notified.

10:45 Angela put new carrier gas online for the GC.
There are now no spare bottles outside.
Purged the line for 2 minutes.

11:25 Valp calls to notify a gain shift for Csi channel
2084. JasonH verifies it and finds the HV setting
is 28V off. We plan to do a full verify of the
calorimeter HVs just before next run.

206 11:57 Ktevana errors two times in a row, ran KFRENDs.

~225 12:24 TeV quench. Estimate for beam:45 min.

13:25 Stops run 14203 to update the CsI HVs.
JasonH and Julie find out that HVs are
off by up to 100V on sdaq !!!! Can't believe
the numbers...

14:30 MCR calls to say beam's gonna ba back soon.
Starts run 14204.

14204 5 14:20 Until spill 5, HV main frame 1 is screwed up.
Julie restore the settings, and then it's fine.

14:50 No histograms for the calorimeter. Looks like
some processes are dead. Stops the run and
do a dart_fresh_start to restart all the
processes.
Run 14204 shouldn't be used.

15:00 Starts run 14205.
Pipeline and ktevana errors.
The weird CsI HVs seeem to be real since the PMT
gain drifts are all outside the plot !!!

15:09 Beam's gone. No estimate

15:15 Reload the HVs from yesterday.
PMT gain drift histo don't improve.
JasonH and Julie are working on it.

========================================================================
Follow up Notes on Previous Shift

CsI 166 : look for comparator problems -> haven't seen any problems.

========================================================================
Items to Follow up on Next Shift


========================================================================

I leave my last shift with scaring problems with CsI HVs...I swear,
it's not my fault ! I didn't screwed up everything before leaving Chicago.

Bye bye everybody,

Sandrine

and Ashkan.

********************************************************************
* LAPLACE Sandrine *
* --------------- *
* Enrico Fermi Institute KTeV 5514,S Blackstone Apt 318*
* 5640, S Ellis ave. at Chicago, IL 60637 *
* Chicago, IL 60637, USA Fermilab USA *
* http://www.geocities.com/Nashville/Rodeo/1190 *
********************************************************************